OpenStudy (anonymous):

Factor 8x^2+14x+3 Show Work!

OpenStudy (anonymous):

= 8x^2 + 12x + 2x +3 = 4x(2x+3) +1(2x+3) = (2x+3)(4x+1)

OpenStudy (anonymous):

@annawilliams40285 have you got it..??

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Do you now how to do this? Solve by factoring 3x^2+5x-2=0

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

Kashan do you know this problem? Solve by factoring 3x^2+5x-2=0

OpenStudy (anonymous):

@Kashan do you know this problem? Solve by factoring 3x^2+5x-2=0

OpenStudy (anonymous):

=3x^2 +6x -x - 2 = 3x(x+2) -1(x+2) = (x+2)(3x-1)

OpenStudy (anonymous):

@Kashan How about this one? Find the discriminant & describe the type of roots 6x^2-2x-4=0

OpenStudy (anonymous):

please post it as a question and send me its link..